Tyler's age is three less than his brother's age. The sum of their ages is 11.
How old is Tyler?

Answer by ewatrrr(24785)

 
Hi,
Let x represent the brother's age ⇒ (x-3) represents Tyler's age
Question States*** The sum of their ages is 11
x + (x-3) = 11
2x - 3 = 11
2x = 14
x = 7, the brother's age. Tyler is 4 years old 7-3+=+4
CHECKING our answer***
7 + 4 = 11
Wish You the Best in your Studies.